OpenStudy (anonymous):

How do I figure out 3 divided by -27 + 2 cubed?

OpenStudy (campbell_st):

well its \[\frac{3}{-27} + 2^3\] using index laws it can be written, after cancelling the 3 as \[\frac{1}{-9} + 2^3 = -3^{-2} + 2^3\]